OpenStudy (anonymous):

y=x*sqrt(2-x^3) , find y'

OpenStudy (alfie):

\[f(x) = y=x*\sqrt{2-x^3}\] Product rule and chain rule. \[1\sqrt{2-x^3}\ + \frac{-3x^2}{2\sqrt{2-x^3}} * x\] So.. \[\frac{2(2-x^3) - 3x^3}{2\sqrt{2-x^3}}\] Which is \[\frac{4-2x^3 - 3x^3}{2\sqrt{2-x^3}}\] Soooo... \[\frac{4 - 5x^3}{2\sqrt{2-x^3}}\]

OpenStudy (anonymous):

Thanks!